OpenStudy (anonymous):

find the slop and y intercept 2y+5x+6=19+5x what would be the answer

OpenStudy (saifoo.khan):

Angel-A

OpenStudy (angela210793):

2y+5x+6=19+5x 2y+5x+6-19-5x=0 2y+6-19=0 2y=13 y=13/2 y intercept (0:13/2) Sai-f? (this would b right?)

OpenStudy (anonymous):

what would b the slope though

OpenStudy (anonymous):

m=0 b=13^2 your slope is 13/2

OpenStudy (anonymous):

firat u have to arrange them at this order: y=mx+b 2y=13 y=13/2